# Model-Agnostic Bounds for Augmented Inverse Probability Weighted Estimators’ Wald-Confidence Interval Coverage in Randomized Controlled Trials

### v1

@@ 6 Discussion
Since $\hat{\sigma}_{k,a}^{2}=P_{n,k}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-\hat{\psi}_{k,a}^{2}$ and ${\mathbb{E}}[\hat{\sigma}_{k,a}^{2}]=P_{*}{\mathbb{E}}_{\hat{Q}_{k,a}}[{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}]-{\mathbb{E}}[\hat{\psi}_{k,a}^{2}]=P_{*}{\mathbb{E}}_{\hat{Q}_{k,a}}[{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}]-\psi_{*,a}^{2}-{\mathrm{Var}}(\hat{\psi}_{k,a})$, by union bound,
$\Pr\left(|\hat{\sigma}_{k,a}^{2}-{\mathbb{E}}[\hat{\sigma}_{k,a}^{2}]|>\frac{n}{|I_{k}|}\sum_{j=1}^{3}\lambda_{j}\right)$
$\leq\Pr\left(|(P_{n,k}-P_{*}){\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}|>\frac{n}{|I_{k}|}\lambda_{1}\right)+\Pr\left(|P_{*}\{{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-{\mathbb{E}}_{\hat{Q}_{k,a}}[{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}]\}|>\frac{n}{|I_{k}|}\lambda_{2}\right)$
$\quad+\Pr\left(|\hat{\psi}_{k,a}^{2}-\psi_{*,a}^{2}-{\mathrm{Var}}(\hat{\psi}_{k,a})|>\frac{n}{|I_{k}|}\lambda_{3}\right).$
@@
Term 1: Let $\zeta:=\sqrt{P_{*}\{{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}\}^{2}}$, $\xi:=P_{*}|{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}|^{3}$, and $\Phi$ denote the cumulative distribution function of standard Gaussian. By Condition 5, $\Pr(\zeta\in[\underline{m},\bar{m}])=\Pr(\xi\leq\bar{m})=1$. By Berry-Esseen Theorem and the exponential tail bound for Gaussian distribution,
$\Pr\left(|(P_{n,k}-P_{*}){\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}|>\frac{n}{|I_{k}|}\lambda_{1}\right)$
$={\mathbb{E}}\left[\Pr\left(\sqrt{|I_{k}|}\left|\frac{(P_{n,k}-P_{*}){\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}}{\zeta}\right|>\sqrt{|I_{k}|}\frac{n}{|I_{k}|\zeta}\lambda_{1}\mid\hat{Q}_{k,a}\right)\right]$
$\leq{\mathbb{E}}\left[2\Phi\left(-\sqrt{|I_{k}|}\frac{n}{|I_{k}|\zeta}\lambda_{1}\right)+{\mathcal{C}}\frac{\xi}{\zeta^{3}\sqrt{|I_{k}|}}\right]$
$\leq{\mathbb{E}}\left[2\exp\left\{-\frac{n^{2}}{2|I_{k}|\zeta^{2}}\lambda_{1}^{2}\right\}+{\mathcal{C}}\frac{\xi}{\zeta^{3}\sqrt{|I_{k}|}}\right]$
@@
${\mathbb{E}}[\{P_{*}({\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-{\mathbb{E}}_{\hat{Q}_{k,a}}[{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}])\}^{2}]\leq 16M^{2}\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}.$
By Chebyshev’s inequality, Term 2 is upper bounded by $\bar{\hat{\mathcal{R}}}(\lambda_{2})$. If Condition S1 holds, then Term 2 is upper bounded by $\bar{\hat{\mathcal{R}}}^{\prime}(\lambda_{2})$.
Term 3: Term 3 is upper bounded by $\Pr\left(|\hat{\psi}_{k,a}^{2}-\psi_{*,a}^{2}|>c\right)$ where $c:=n\lambda_{3}/|I_{k}|-{\mathrm{Var}}(\hat{\psi}_{k,a})>0$. By Lemma S4, this probability is further upper bounded by
$\Pr\left(|\hat{\psi}_{k,a}-\psi_{*,a}|>(\sqrt{2}-1)\left\{\sqrt{c}\wedge\frac{c}{|\psi_{*,a}|}\right\}\right),$
@@
$={\mathbb{E}}\left[\Pr\left(\sqrt{|I_{k}|}\frac{\left|P_{n,k}D_{a}(\hat{Q}_{k,a},\psi_{*,a})\right|}{\sqrt{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}}}>(\sqrt{2}-1)\frac{\sqrt{|I_{k}|}}{\sqrt{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}}}\left\{\sqrt{c}\wedge\frac{c}{|\psi_{*,a}|}\right\}\mid\hat{Q}_{k,a}\right)\right]$
$\leq{\mathbb{E}}\left[2\Phi\left(-(\sqrt{2}-1)\frac{\sqrt{|I_{k}|}}{\sqrt{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}}}\left\{\sqrt{c}\wedge\frac{c}{|\psi_{*,a}|}\right\}\right)+{\mathcal{C}}\frac{P_{*}|D_{a}(\hat{Q}_{k,a},\psi_{*,a})|^{3}}{\left\{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}\right\}^{3/2}\sqrt{|I_{k}|}}\right]$
$\leq{\mathbb{E}}\left[2\exp\left\{-\frac{(\sqrt{2}-1)^{2}}{2}\frac{|I_{k}|}{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}}\left(c\wedge\frac{c^{2}}{|\psi_{*,a}|^{2}}\right)\right\}+{\mathcal{C}}\frac{P_{*}|D_{a}(\hat{Q}_{k,a},\psi_{*,a})|^{3}}{\left\{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}\right\}^{3/2}\sqrt{|I_{k}|}}\right]$
@@
$=\sum_{k=1}^{K}\frac{|I_{k}|}{n}\left\{{\mathbb{E}}[\hat{\sigma}_{k,a}^{2}]-\sigma_{\#,a}^{2}\right\}$
$=\sum_{k=1}^{K}\frac{|I_{k}|}{n}\left\{{\mathbb{E}}[P_{n,k}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-\hat{\psi}_{k,a}^{2}]-P_{*}{\mathcal{T}}_{a}(Q_{\#,a})^{2}+\psi_{*,a}^{2}\right\}$
$=\sum_{k=1}^{K}\frac{|I_{k}|}{n}\Bigg\{{\mathbb{E}}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}\right]+2{\mathbb{E}}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})(Q_{\#,a}-Q_{*,a})\right]$
@@
Under Condition 3, by the triangle inequality, Cauchy-Schwarz inequality, and Lemma S8, $|\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}|\leq\sum_{k=1}^{K}\hat{S}_{k}$. Under Condition 8, $|\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}|\leq\sum_{k=1}^{K}\hat{S}_{k}^{\prime}$ under Condition 3 by Lemma S8. The bounds for $|\sigma_{\dagger,a}-\sigma_{\#,a}|$ follow from the fact that $|\sigma_{\dagger,a}-\sigma_{\#,a}|=|\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}|/(\sigma_{\dagger,a}+\sigma_{\#,a})\leq|\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}|/\sigma_{\#,a}$, where $\sigma_{\#,a}\geq\sigma_{*,a}>0$ by Condition 1 and the fact that $\sigma_{*,a}^{2}$ is the efficiency bound. ∎
@@
$\Pr\left(|\sqrt{n}\{\hat{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}-z_{\alpha}(\hat{\sigma}_{a}-\sigma_{\dagger,a})|>\bar{\eta}\right)$
@@
$\hat{\hat{\mathcal{R}}}(\eta,\lambda_{2})$ $:=\frac{|I_{k}|}{n}\frac{1-\tau_{\pi}}{\tau_{\pi}\eta^{2}}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}+{\mathbbm{1}}(z_{\alpha}\neq 0)\bar{\hat{\mathcal{R}}}\left(\frac{\sigma_{\dagger,a}}{|z_{\alpha}|}\lambda_{2}\right).$
Additionally under Conditions 7 and S1, for any constant $q>0$ and all $k=1,\ldots,K$, $\hat{\hat{\mathcal{R}}}(\eta_{k},\lambda_{k,2})$ can be replaced by $\hat{\hat{\mathcal{R}}}^{\prime}(\eta_{k},\lambda_{k,2},q)$, which is defined pointwise as
$\hat{\hat{\mathcal{R}}}^{\prime}(\eta,\lambda_{2},q)$
$:=2\exp\left\{-\frac{n\eta^{2}}{2|I_{k}|\hat{s}_{1}^{2}\left[{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}+\hat{s}_{2}\sqrt{2q{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\log n}\right]^{2}}\right\}$
$\qquad+2n^{-q}+{\mathbbm{1}}(z_{\alpha}\neq 0)\bar{\hat{\mathcal{R}}}^{\prime}\left(\frac{\sigma_{\dagger,a}}{|z_{\alpha}|}\lambda_{2}\right).$
@@
$\Pr\left(|\sqrt{n}\{\hat{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}-z_{\alpha}(\hat{\sigma}_{a}-\sigma_{\dagger,a})|>\sum_{k=1}^{K}\eta_{k}+\sum_{k=1}^{K}\sum_{j=1}^{3}\lambda_{k,j}\right)$
$\leq\sum_{k=1}^{K}\Pr\left(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta_{k}\right)+{\mathbbm{1}}(z_{\alpha}\neq 0)\Pr\left(|\hat{\sigma}_{a}-\sigma_{\dagger,a}|>\sum_{k=1}^{K}\sum_{j=1}^{3}\lambda_{k,j}/|z_{\alpha}|\right).$
Here, $\Pr\left(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta_{k}\right)$ can be bounded as in Lemma S7. Since $\sigma_{\dagger,a}+\hat{\sigma}_{a}\geq\sigma_{\dagger,a}>0$, if $z_{\alpha}\neq 0$,
$\Pr\left(|\hat{\sigma}_{a}-\sigma_{\dagger,a}|>\sum_{k=1}^{K}\sum_{j=1}^{3}\lambda_{k,j}/|z_{\alpha}|\right)$
$\leq\Pr\left(|\hat{\sigma}_{a}^{2}-\sigma_{\dagger,a}^{2}|>\sum_{k=1}^{K}\sum_{j=1}^{3}\lambda_{k,j}\sigma_{\dagger,a}/|z_{\alpha}|\right)$
$=\Pr\left(\left|\frac{1}{n}\sum_{k=1}^{K}|I_{k}|(\hat{\sigma}_{k,a}^{2}-{\mathbb{E}}[\hat{\sigma}_{k,a}^{2}])\right|>\sum_{k=1}^{K}\sum_{j=1}^{3}\lambda_{k,j}\sigma_{\dagger,a}/|z_{\alpha}|\right)$
$\leq\sum_{k=1}^{K}\Pr\left(|\hat{\sigma}_{k,a}^{2}-{\mathbb{E}}[\hat{\sigma}_{k,a}^{2}]|>\frac{n}{|I_{k}|}\sum_{j=1}^{3}\frac{\lambda_{k,j}\sigma_{\dagger,a}}{|z_{\alpha}|}\right),$
@@
$\left|\Pr(\sqrt{n}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\hat{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
@@
$\eta_{k}=\left\{\frac{|I_{k}|}{n}\frac{1-\tau_{\pi}}{\tau_{\pi}}\sigma_{\#,a}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\right\}^{1/3},\quad\lambda_{k,1}=\begin{cases}\frac{|z_{\alpha}|\sqrt{2q|I_{k}|\bar{m}\log n}}{n\sigma_{\dagger,a}}&(\text{if }z_{\alpha}\neq 0)\\ n^{-q}&(\text{if }z_{\alpha}=0)\end{cases},$
$\lambda_{k,2}=\left\{16M^{2}\frac{|I_{k}|^{2}z_{\alpha}^{2}\sigma_{\#,a}}{n^{2}\sigma_{\dagger,a}^{2}}\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\right\}^{1/3},$

### v2

@@ 6 Discussion
Since $\hat{\sigma}_{k,a}^{2}=P_{n,k}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-\hat{\psi}_{k,a}^{2}$ and ${E}[\hat{\sigma}_{k,a}^{2}]=P_{*}{E}_{\hat{Q}_{k,a}}[{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}]-{E}[\hat{\psi}_{k,a}^{2}]=P_{*}{E}_{\hat{Q}_{k,a}}[{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}]-\psi_{*,a}^{2}-{\mathrm{var}}(\hat{\psi}_{k,a})$, by union bound,
${\mathrm{pr}}\left(|\hat{\sigma}_{k,a}^{2}-{E}[\hat{\sigma}_{k,a}^{2}]|>\frac{n}{|I_{k}|}\sum_{j=1}^{3}\lambda_{j}\right)$
$\leq{\mathrm{pr}}\left(|(P_{n,k}-P_{*}){\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}|>\frac{n}{|I_{k}|}\lambda_{1}\right)+{\mathrm{pr}}\left(|P_{*}\{{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-{E}_{\hat{Q}_{k,a}}[{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}]\}|>\frac{n}{|I_{k}|}\lambda_{2}\right)$
$\quad+{\mathrm{pr}}\left(|\hat{\psi}_{k,a}^{2}-\psi_{*,a}^{2}-{\mathrm{var}}(\hat{\psi}_{k,a})|>\frac{n}{|I_{k}|}\lambda_{3}\right).$
@@
Term 1: Let $\zeta:=\sqrt{P_{*}\{{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}\}^{2}}$, $\xi:=P_{*}|{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}|^{3}$, and $\Phi$ denote the cumulative distribution function of standard Gaussian. By Condition 5, ${\mathrm{pr}}(\zeta\in[\underline{m},\bar{m}])={\mathrm{pr}}(\xi\leq\bar{m})=1$. By Berry-Esseen Theorem and the exponential tail bound for Gaussian distribution,
${\mathrm{pr}}\left(|(P_{n,k}-P_{*}){\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}|>\frac{n}{|I_{k}|}\lambda_{1}\right)$
$={E}\left[{\mathrm{pr}}\left(\sqrt{|I_{k}|}\left|\frac{(P_{n,k}-P_{*}){\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}}{\zeta}\right|>\sqrt{|I_{k}|}\frac{n}{|I_{k}|\zeta}\lambda_{1}\mid\hat{Q}_{k,a}\right)\right]$
$\leq{E}\left[2\Phi\left(-\sqrt{|I_{k}|}\frac{n}{|I_{k}|\zeta}\lambda_{1}\right)+{\mathcal{C}}\frac{\xi}{\zeta^{3}\sqrt{|I_{k}|}}\right]$
$\leq{E}\left[2\exp\left\{-\frac{n^{2}}{2|I_{k}|\zeta^{2}}\lambda_{1}^{2}\right\}+{\mathcal{C}}\frac{\xi}{\zeta^{3}\sqrt{|I_{k}|}}\right]$
@@
${E}[\{P_{*}({\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-{E}_{\hat{Q}_{k,a}}[{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}])\}^{2}]\leq 16M^{2}\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}.$
By Chebyshev’s inequality, Term 2 is upper bounded by $\bar{\hat{\mathcal{R}}}(\lambda_{2})$. If Condition 8 holds, then Term 2 is upper bounded by $\bar{\hat{\mathcal{R}}}^{\prime}(\lambda_{2})$.
Term 3: Term 3 is upper bounded by ${\mathrm{pr}}\left(|\hat{\psi}_{k,a}^{2}-\psi_{*,a}^{2}|>c\right)$ where $c:=n\lambda_{3}/|I_{k}|-{\mathrm{var}}(\hat{\psi}_{k,a})>0$. By Lemma S4, this probability is further upper bounded by
${\mathrm{pr}}\left(|\hat{\psi}_{k,a}-\psi_{*,a}|>(\sqrt{2}-1)\left\{\sqrt{c}\wedge\frac{c}{|\psi_{*,a}|}\right\}\right),$
@@
$={E}\left[{\mathrm{pr}}\left(\sqrt{|I_{k}|}\frac{\left|P_{n,k}D_{a}(\hat{Q}_{k,a},\psi_{*,a})\right|}{\sqrt{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}}}>(\sqrt{2}-1)\frac{\sqrt{|I_{k}|}}{\sqrt{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}}}\left\{\sqrt{c}\wedge\frac{c}{|\psi_{*,a}|}\right\}\mid\hat{Q}_{k,a}\right)\right]$
$\leq{E}\left[2\Phi\left(-(\sqrt{2}-1)\frac{\sqrt{|I_{k}|}}{\sqrt{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}}}\left\{\sqrt{c}\wedge\frac{c}{|\psi_{*,a}|}\right\}\right)+{\mathcal{C}}\frac{P_{*}|D_{a}(\hat{Q}_{k,a},\psi_{*,a})|^{3}}{\left\{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}\right\}^{3/2}\sqrt{|I_{k}|}}\right]$
$\leq{E}\left[2\exp\left\{-\frac{(\sqrt{2}-1)^{2}}{2}\frac{|I_{k}|}{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}}\left(c\wedge\frac{c^{2}}{|\psi_{*,a}|^{2}}\right)\right\}+{\mathcal{C}}\frac{P_{*}|D_{a}(\hat{Q}_{k,a},\psi_{*,a})|^{3}}{\left\{P_{*}D_{a}(\hat{Q}_{k,a},\psi_{*,a})^{2}\right\}^{3/2}\sqrt{|I_{k}|}}\right]$
@@
$=\sum_{k=1}^{K}\frac{|I_{k}|}{n}\left\{{E}[\hat{\sigma}_{k,a}^{2}]-\sigma_{\#,a}^{2}\right\}$
$=\sum_{k=1}^{K}\frac{|I_{k}|}{n}\left\{{E}[P_{n,k}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-\hat{\psi}_{k,a}^{2}]-P_{*}{\mathcal{T}}_{a}(Q_{\#,a})^{2}+\psi_{*,a}^{2}\right\}$
$=\sum_{k=1}^{K}\frac{|I_{k}|}{n}\Bigg\{{E}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}\right]+2{E}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})(Q_{\#,a}-Q_{*,a})\right]$
@@
Under Condition 3, by the triangle inequality, Cauchy-Schwarz inequality, and Lemma S8, $|\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}|\leq\sum_{k=1}^{K}\hat{S}_{k}$. Under Condition 9, $|\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}|\leq\sum_{k=1}^{K}\hat{S}_{k}^{\prime}$ under Condition 3 by Lemma S8. The bounds for $|\sigma_{\dagger,a}-\sigma_{\#,a}|$ follow from the fact that $|\sigma_{\dagger,a}-\sigma_{\#,a}|=|\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}|/(\sigma_{\dagger,a}+\sigma_{\#,a})\leq|\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}|/\sigma_{\#,a}$, where $\sigma_{\#,a}\geq\sigma_{*,a}>0$ by Condition 1 and the fact that $\sigma_{*,a}^{2}$ is the efficiency bound. ∎
@@
${\mathrm{pr}}\left(|\sqrt{n}\{\hat{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}-z_{\alpha}(\hat{\sigma}_{a}-\sigma_{\dagger,a})|>\bar{\eta}\right)$
@@
$\hat{\hat{\mathcal{R}}}(\eta,\lambda_{2})$ $:=\frac{|I_{k}|}{n}\frac{1-\tau_{\pi}}{\tau_{\pi}\eta^{2}}{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}+{\mathbbm{1}}(z_{\alpha}\neq 0)\bar{\hat{\mathcal{R}}}\left(\frac{\sigma_{\dagger,a}}{|z_{\alpha}|}\lambda_{2}\right).$
Additionally under Conditions 7 and 8, for any constant $q>0$ and all $k=1,\ldots,K$, $\hat{\hat{\mathcal{R}}}(\eta_{k},\lambda_{k,2})$ can be replaced by $\hat{\hat{\mathcal{R}}}^{\prime}(\eta_{k},\lambda_{k,2},q)$, which is defined pointwise as
$\hat{\hat{\mathcal{R}}}^{\prime}(\eta,\lambda_{2},q)$ $:=2\exp(-\eta^{\prime})+2\hat{a}_{1}n^{-q}+2\hat{b}_{1}\sqrt{\log n/n}+{\mathbbm{1}}(z_{\alpha}\neq 0)\bar{\hat{\mathcal{R}}}^{\prime}\left(\frac{\sigma_{\dagger,a}}{|z_{\alpha}|}\lambda_{2}\right).$
where $\eta^{\prime}=\eta^{\prime}(\eta)$ is defined in Lemma S7.
@@
${\mathrm{pr}}\left(|\sqrt{n}\{\hat{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}-z_{\alpha}(\hat{\sigma}_{a}-\sigma_{\dagger,a})|>\sum_{k=1}^{K}\eta_{k}+\sum_{k=1}^{K}\sum_{j=1}^{3}\lambda_{k,j}\right)$
$\leq\sum_{k=1}^{K}{\mathrm{pr}}\left(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta_{k}\right)+{\mathbbm{1}}(z_{\alpha}\neq 0){\mathrm{pr}}\left(|\hat{\sigma}_{a}-\sigma_{\dagger,a}|>\sum_{k=1}^{K}\sum_{j=1}^{3}\lambda_{k,j}/|z_{\alpha}|\right).$
Here, ${\mathrm{pr}}\left(|I_{k}||P_{n,k}H_{a}(\hat{Q}_{k,a})|>n^{1/2}\eta_{k}\right)$ can be bounded as in Lemma S7. Since $\sigma_{\dagger,a}+\hat{\sigma}_{a}\geq\sigma_{\dagger,a}>0$, if $z_{\alpha}\neq 0$,
${\mathrm{pr}}\left(|\hat{\sigma}_{a}-\sigma_{\dagger,a}|>\sum_{k=1}^{K}\sum_{j=1}^{3}\lambda_{k,j}/|z_{\alpha}|\right)$
$\leq{\mathrm{pr}}\left(|\hat{\sigma}_{a}^{2}-\sigma_{\dagger,a}^{2}|>\sum_{k=1}^{K}\sum_{j=1}^{3}\lambda_{k,j}\sigma_{\dagger,a}/|z_{\alpha}|\right)$
$={\mathrm{pr}}\left(\left|\frac{1}{n}\sum_{k=1}^{K}|I_{k}|(\hat{\sigma}_{k,a}^{2}-{E}[\hat{\sigma}_{k,a}^{2}])\right|>\sum_{k=1}^{K}\sum_{j=1}^{3}\lambda_{k,j}\sigma_{\dagger,a}/|z_{\alpha}|\right)$
$\leq\sum_{k=1}^{K}{\mathrm{pr}}\left(|\hat{\sigma}_{k,a}^{2}-{E}[\hat{\sigma}_{k,a}^{2}]|>\frac{n}{|I_{k}|}\sum_{j=1}^{3}\frac{\lambda_{k,j}\sigma_{\dagger,a}}{|z_{\alpha}|}\right),$
@@
$\left|{\mathrm{pr}}(\sqrt{n}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\hat{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
@@
$\eta_{k}=\left\{\frac{|I_{k}|}{n}\frac{1-\tau_{\pi}}{\tau_{\pi}}\sigma_{\#,a}{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\right\}^{1/3},\quad\lambda_{k,1}=\begin{cases}\frac{|z_{\alpha}|\sqrt{2q|I_{k}|\bar{m}\log n}}{n\sigma_{\dagger,a}}&(\text{if }z_{\alpha}\neq 0)\\ n^{-q}&(\text{if }z_{\alpha}=0)\end{cases},$
$\lambda_{k,2}=\left\{16M^{2}\frac{|I_{k}|^{2}z_{\alpha}^{2}\sigma_{\#,a}}{n^{2}\sigma_{\dagger,a}^{2}}\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\right\}^{1/3},$